# Mapping the Potential of Explainable Artificial Intelligence (XAI) for Fairness Along the AI Lifecycle

### v2

@@ 3.6. Empowering Fairness
“From the perspective of individuals affected by automated decision-making, we propose three aims for explanations: (1) to inform and help the individual understand why a particular decision was reached, (2) to provide grounds to contest the decision if the outcome is undesired, and (3) to understand what could be changed to receive a desired result in the future, based on the current decision-making model.” (Wachter et al., 2017, p. 2)
@@
Fairness considerations do not end after an AI model has been designed and decisions have been made. Empowering fairness refers to the ability of affected parties to take effective actions regarding their outcome of a particular decision, e.g., by contesting decisions or seeking recourse.
@@
In their article on the right to explanation, Vredenburgh (2022) proposes two types of self-informed advocacy: retrospectively, affected parties should be able to identify the accountable entity to demand remedy for unfair treatment (viz., responsibility); prospectively, contesting and recourse options should enable affected parties to actively improve upon their possibly unfair outcome (viz., agency). Our conception of empowering fairness strongly relates to Vredenburgh’s forward-looking self-informed advocacy. It more generally relates to attempts of conceptualizing what makes a fair explanation in the context of the right to explanation (Hacker and Passoth, 2022).
@@
The Role of XAI. XAI could be useful for empowering fairness because it may help acquire the information to be communicated. Counterfactual explanations have been claimed to provide a promising solution to inform and empower affected parties (Wachter et al., 2017; Asher et al., 2022). For example, they can clarify what combination of feature values would lead to a different outcome. Therefore, counterfactual explanations can provide valuable information on recourse, e.g., by telling a loan applicant how much more she would need to earn in order to be granted a loan (Wachter et al., 2017). While research on counterfactual explanations is growing rapidly, they do not come without limitations regarding their validity (Guidotti, 2022), underlying assumptions (Hu and Kohler-Hausmann, 2020), or susceptibility to manipulations (Slack et al., 2021). Extending the scope beyond model-centered explanations, XAI might also involve practical information about responsible contact persons, guidance on how to seek redress, or collaborative platforms for affected parties to share experienced outcomes (Le Merrer and Trédan, 2020).
3.7. Long-Term Fairness
“Accuracy, discrimination, and security characteristics of a system can change over time as well. Simply testing for these problems at training time […] is not adequate for high-stakes, human-centered, or regulated ML systems. Accuracy, discrimination, and security should be monitored in real-time and over time, as long as a model is deployed.” (Gill et al., 2020, p. 18)
@@ 3.7. Long-Term Fairness
The long-term impact of an AI model can be analyzed from several perspectives. Arif Khan et al. (2022) contrast “formal” and “substantive” equality of opportunity where a forward-facing view of fair life chances also accounts for affected parties’ future prospects of success (as opposed to ensuring fair contests at a discrete point in time). Since such conceptions make assumptions about structural disadvantages and future prospects of affected parties, they require much broader and longitudinal evaluation. Further, due to concept drift, a model’s formal fairness properties can change over time and should be monitored.
Stage of the Lifecycle. We map long-term fairness primarily onto usage & inference, noting that it may encompass all future iterations of the AI lifecycle until the AI system is shut down.
The Role of XAI. Monitoring tools may help to track changes of fairness metrics and identify situations where interventions are necessary (Castelnovo et al., 2021; Panigutti et al., 2021; Ghosh et al., 2022). Another long-term fairness impact is strategic gaming behavior that arises from transparent models (Lepri et al., 2018). For example, loan applicants who receive counterfactual explanations exactly describing how to reach the decision threshold are prone to create a game-theoretic situation where information itself can be unfairly distributed among clients. We suspect several other forms of long-term fairness issues to emerge that are currently under-explored in the literature. For example, Liu et al. (2018) model the impact of formal fairness on the underlying population over time, which Hardt et al. (2022) coined as “performative power”. Novel forms of XAI may help to anticipate and evaluate such dynamics.
3.8. Informational Fairness
Transparency […] is valuable because and in so far as it enables the individuals, who are subjected to algorithmic decision-making, to assess whether these decisions are morally and politically justifiable. (Loi et al., 2021, p. 254)
@@ 3.8. Informational Fairness
Informational fairness was originally introduced as a psychological construct in the context of organizational justice (Colquitt, 2001) to test whether the communication accompanying a decision is candid, truthful, reasonable, timely, and specific. Our conception of informational fairness is inspired by this construct and corresponds to a great extent to Loi et al. (2021)’s concept of design publicity. In line with our idea of broad XAI, Loi et al. (2021) demand a form of transparency that not only explains a model and its underlying functioning but also the goals and values that went into its design and how these are embedded in the model. A similar distinction is made by Walmsley (2021) who differentiates between functional transparency concerned with the inner workings of an AI model and outward transparency related to the communication with stakeholders.
@@
The Role of XAI. In the case of informational fairness, XAI is not an aid to but the desideratum itself. Following this idea, we provide examples of what could be made transparent and communicated to affected parties for each fairness desideratum. Regarding fairness understanding, developers and deployers could justify their fairness objectives, delineate potential tradeoffs, and elucidate the mechanisms through which AI aims to achieve them (Loi et al., 2021). For data fairness, they could explain how the training data is composed highlighting potential sources of bias (Anik and Bunt, 2021). Explaining the training objectives and outlining potential uncertainties or goal conflicts might be appropriate to address formal fairness (Mitchell et al., 2019). Regarding perceived fairness, explaining the functions and limitations of XAI-generated information could enhance understanding to approach “appropriate fairness perceptions” (Schoeffer and Kuehl, 2021). Transparency about the role of AI and the responsibilities of a human decision-maker within the decision-making context might be desirable for fairness with human oversight (Martin, 2019). The idea of empowering fairness relies on truthful, understandable, and actionable information (Watson and Floridi, 2021). Lastly, regarding long-term fairness, unexpected downstream impacts and the factors driving potentially unfair dynamics can be communicated to stakeholders (Slack et al., 2020).
4. The COMPAS Case
@@
To illustrate the utility of our mapping, we describe how XAI could help address fairness throughout the lifecycle of a high-stake AI system. As an example, we consider the recidivism prediction software COMPAS developed by Northpointe (today rebranded to equivant Supervision). COMPAS has been installed in many US-American jurisdictions in order to predict whether a defendant will likely commit another crime in the near future. Judges use this information, e.g., for decisions about who they release on bail. However, COMPAS has been criticized by investigative journalists at ProPublica for disadvantaging Black people (Angwin et al., 2022). This is especially problematic given the history of systematic discrimination and marginalization of Black people in the US (Pettit and Gutierrez, 2018). In this section, we illustrate both how Northpointe could have addressed different fairness desiderata during system development and how they could still, given COMPAS’ continued use, address some of them. Of course, XAI is not the only means to this end and is not to be conceived as an ethical panacea. Notably, our illustration also presupposes an inherent motivation to actually address fairness desiderata, which is not necessarily given in the context of a profit-oriented proprietary system like COMPAS.
@@
Let us proceed along the AI lifecycle. First, Northpointe could have used XAI to ensure that the development is based on an appropriate fairness understanding. ProPublica’s analysis of COMPAS shows that although it was tested for some fairness objective (viz., predictive parity), it falls short on other fairness objectives (viz., equalized odds) in problematic ways (Chouldechova, 2017). Thus, predictive parity may have been an inadequate fairness objective. Northpointe could also have based their development on insights about the predictive value and correlations of certain features. It has been shown that defendants’ age and number of previous crimes are most predictive for recidivism (Dressel and Farid, 2018; Rudin, 2019) but also highly correlated to race. Based on such insights, Northpointe could have taken a position on how to treat such features considering pre-existing systematic discrimination.
@@
To address data fairness, Northpointe would have needed to ensure that the data set adequately represents demographic groups targeted by the system and to be aware of existing structural relationships such as statistically higher crime rates in predominantly Black neighborhoods (Binns, 2018). Descriptive statistics could have already pointed them towards biases introduced in the data collection process—a reason to reiterate the data collection stage to apply new data collection strategies (Li et al., 2022). Statistical analysis could also have helped unveil traces of systematic discrimination in the data (e.g. that Black people are more likely to be arrested for minor offenses due to increased police presence in Black neighborhoods). Northpointe could have used such insights either to change their data collection strategies (e.g., collecting features that are less correlated to race), or they could have used them during the feature selection phase (e.g., to mitigate systematic discrimination at the data-level (Pradhan et al., 2022)).
@@
Regarding formal fairness, Northpointe tested COMPAS for equality of error rates (viz., predictive parity). By contrast, journalists at ProPublica tested for the balance of true and false positives (viz., equalized odds) (Chouldechova, 2017). Selecting appropriate formal fairness metrics is an intricate endeavor (De-Arteaga et al., 2022). However, had Northpointe tested COMPAS for equalized odds, too, they might have prevented substantial societal harms. Today, developers can draw on an extensive suite of fairness testing tools for a comprehensive formal fairness assessment (Linardatos et al., 2020; Chen et al., 2023; Deck et al., 2023).
@@
Northpointe could have also detected flaws by engaging with stakeholders and gathering feedback on perceived fairness. Outcome statistics, fairness metrics, or model explanations (e.g., in the form of counterfactuals) could have been presented to a focus group of diverse stakeholders to assess and discuss fairness perceptions (cf. (Stumpf et al., 2021)). Potential concerns could be handled by reconciling the different perspectives, and even today (after deployment) Northpointe could continue to monitor fairness perceptions by providing feedback channels.
@@
Our discussion thus far has primarily focused on measures that COMPAS’ developers could have taken early in the AI lifecycle. But even now, during COMPAS’ continued usage, Northpointe could still address several fairness issues via XAI. First, they could seek fairness with human oversight. This is a particularly relevant desideratum because COMPAS does not make autonomous decisions but informs the decisions of judges. So, Northpointe could familiarize judges with the basic functioning, strengths, limitations, and uncertainties of the system as well as the factors driving the risk score to ensure judges will use it appropriately. Further, analyzing or supervising the judges’ interactions with the system and its explanations might help to achieve fairness objectives more effectively (Teodorescu et al., 2021).
@@
Northpointe could also foster empowering fairness by providing defendants with two kinds of explanations: First, global explanations (Dodge et al., 2019) about the general functioning of COMPAS could be communicated. Where relevant factors are under a person’s control (as, e.g., with defendants’ history of misdemeanor or substance abuse), insights about their impact can allow people to make life choices decreasing their risk scores. Second, local explanations (Dodge et al., 2019) could provide information about an individual’s risk score granting the opportunity to (justifiably) contest the score, e.g., by denouncing discriminatory treatment (John-Mathews, 2022), or to (effectively) seek recourse, e.g., by signaling plans to go into rehab (Wachter et al., 2017).
@@
Moreover, Northpointe could monitor long-term fairness by tracking the specified formal fairness objective(s) over time and accordingly adjust the model over several iterations of the AI lifecycle (Castelnovo et al., 2021). Further, changes in defendants’ behavior due to specified fairness objectives or increased transparency could be examined. For example, precise global or counterfactual explanations might allow defendants to strategically game the system in unexpected ways (Lepri et al., 2018), which might—similar to unfair recourse (Gupta et al., 2019)—be easier for some than for others.
@@
Finally, Northpointe could have aimed, and still could aim, to ensure informational fairness. To this end, they would have needed to document fairness-related information and communicate it accordingly. For example, Northpointe could have used model (fairness) cards similar to (Mitchell et al., 2019), which would have clarified many of the issues that were only revealed due to the investigative work of Angwin et al. (2022). Most importantly perhaps, Northpointe could have been candid about the underlying business and fairness objectives from the start, ideally making transparent decisions about tradeoffs (e.g., when looking at predictive parity rather than equalized odds) and justifying why an AI system was an appropriate tool for criminal risk prediction in the first place. In communication with affected parties, COMPAS could be complemented with truthful, understandable, and relevant information (Asher et al., 2022) explaining the underlying logic. At the same time, jurisdictions could be more transparent about how exactly COMPAS is employed and how judges interact with the system. Overall, the strategic use of XAI along the AI lifecycle could benefit all fairness desiderata regarding a model like COMPAS—given sincere intentions to actually pursue these desiderata.
5. Conclusion and Outlook
@@
We distilled eight fairness desiderata from interdisciplinary literature, mapped them onto the AI lifecycle, and discussed how XAI might be leveraged to contribute to fulfilling them. Finally, we illustrated the utility of our approach by applying it to the COMPAS case. The overall picture we paint (see Figure 2) highlights that to effectively design and utilize XAI for fairness, it is paramount to reflect about which fairness desideratum one seeks to fulfill. Before closing, we shall briefly comment on some limitations of our proposal and point to avenues for future work.
@@
Generalizability. Our discussion has focused primarily on high-risk applications and more traditional decision-support systems. To what extent does our proposal generalize to low-risk applications? And what about generative AI? Consider Google’s multi-modal AI-chatbot Gemini, which has recently received bad publicity for its inaccurate historical depictions—which were driven by misdirected fairness interventions (The, 2024). We suspect that the usage of (seemingly low-risk) AI systems at scale will accumulate to significant societal impacts bearing more subtle threats to fairness such as, e.g., representational harms (Barocas et al., 2017) in the Gemini case. Thus, although contemporary regulation (like the European AI Act) focuses predominantly on high-risk systems, fairness is an important concern for low-risk applications, too. Beyond that, the Gemini case illustrates that fairness considerations affect generative AI just as much as more traditional AI systems for decision-making. Given the rapid advancement and broad adoption of generative AI, we believe that fairness considerations will be indispensable in this context, though they only start to gain traction (Friedrich et al., 2023; Neumann et al., 2024). What makes them particularly challenging is that traditional stakeholder roles may no longer hold: in the Gemini case, users and affected parties are the same people; and ChatGPT users creating personalized chatbots are no longer “just users” either. Thus, to ensure fairness, future research might not only have to refine fairness desiderata but also rethink stakeholder categories.
Actionability. We acknowledge that our proposal does not present an actionable process model for fair software engineering; nor does it provide regulatory guidelines to be enforced by legal institutions. Still, it may serve as a useful roadmap for researchers, developers, and regulators. For researchers, it provides a starting point for a truly interdisciplinary discourse going beyond discussions of algorithmic fairness focusing on technical aspects of AI systems. For developers, our mapping provides urgently needed guidance as to what fairness challenges should be addressed when developing specific guidelines and requirements (see Hacker (2022)). More specifically, it may help determine under which circumstances human oversight is beneficial (Teodorescu et al., 2021; Schoeffer et al., 2024) or how responsibilities can be attributed along the AI lifecycle (Raji et al., 2020). Naturally, the implications of our approach will vary with the precise settings; so fine-tuning may be needed for different business, societal, and legal contexts.

### v4

@@ 3.6 Empowering Fairness
“From the perspective of individuals affected by automated decision-making, we propose three aims for explanations: (1) to inform and help the individual understand why a particular decision was reached, (2) to provide grounds to contest the decision if the outcome is undesired, and (3) to understand what could be changed to receive a desired result in the future, based on the current decision-making model.” [79, p. 2]
@@
Fairness considerations do not end after an AI model has been designed and decisions have been made. Empowering fairness refers to the ability of affected parties to take effective actions regarding the outcome of a particular decision, e.g., by contesting decisions or seeking recourse.
@@
In her article on the right to explanation, Vredenburgh [77] proposes two types of self-informed advocacy: retrospectively, affected parties should be able to identify the accountable entity to demand remedy for unfair treatment (viz., responsibility); prospectively, contesting and recourse options should enable affected parties to actively improve upon their possibly unfair outcome (viz., agency). Our conception of empowering fairness strongly relates to Vredenburgh’s forward-looking self-informed advocacy. It more generally relates to attempts to conceptualize what makes a fair explanation in the context of the right to explanation [87].
@@
The Role of XAI. XAI could be useful for empowering fairness because it may help acquire the information to be communicated. Contrastive explanations (answering the question “Why P rather than Q?”) are commonly proposed to provide intuitive entry points to engage with affected parties [79, 127]. For instance, counterfactual explanations have been claimed to provide a promising solution to inform and empower affected parties by clarifying what combination of feature values would lead to a different outcome. [79, 80]. Therefore, they can provide valuable information required for algorithmic recourse, e.g., by providing actionable recommendations to a loan applicant on what to do to be granted a loan in the future [128, 129]. While research on counterfactual explanations is growing rapidly, they do not come without limitations regarding their actionability [128], validity [130], underlying assumptions [131], or susceptibility to manipulations [132]. Extending the scope beyond model-centered explanations, XAI might also involve practical information about responsible contact persons, guidance on how to seek redress, or collaborative platforms for affected parties to share experienced outcomes [117].
3.7 Long-Term Fairness
“Accuracy, discrimination, and security characteristics of a system can change over time as well. Simply testing for these problems at training time […] is not adequate for high-stakes, human-centered, or regulated ML systems. Accuracy, discrimination, and security should be monitored in real-time and over time, as long as a model is deployed.” [133, p. 18]
@@ 3.7 Long-Term Fairness
The long-term impact of an AI model can be analyzed from several perspectives. Arif Khan et al. [134] contrast “formal” and “substantive” equality of opportunity where a forward-facing view of fair life chances also accounts for affected parties’ future prospects of success (as opposed to ensuring fair contests at a discrete point in time). Since such conceptions make assumptions about structural disadvantages and future prospects of affected parties, they require much broader and longitudinal evaluation. Further, due to concept drift [135], a model’s formal fairness properties can change over time and should be monitored.
Stage of the Lifecycle. We map long-term fairness primarily onto inference & usage, noting that it may encompass all future iterations of the AI lifecycle until the AI system is shut down.
The Role of XAI. Monitoring tools may help to track changes in fairness metrics and identify situations where interventions are necessary [83, 85, 82]. Another long-term fairness impact is strategic gaming behavior that arises from transparent models [136]. For example, loan applicants who receive counterfactual explanations exactly describing how to reach the decision threshold are prone to create a game-theoretic situation where information itself can be unfairly distributed among clients. We suspect several other forms of long-term fairness issues will emerge that are currently under-explored in the literature. For example, Liu et al. [137] model the impact of formal fairness on the underlying population over time, which Hardt et al. [138] coined as “performative power”. Novel forms of XAI may help to anticipate and evaluate such dynamics.
3.8 Informational Fairness
Transparency […] is valuable because and in so far as it enables the individuals, who are subjected to algorithmic decision-making, to assess whether these decisions are morally and politically justifiable. [52, p. 254]
@@ 3.8 Informational Fairness
Informational fairness was originally introduced as a psychological construct in the context of organizational justice [70] to test whether the communication accompanying a decision is candid, truthful, reasonable, timely, and specific. Our conception of informational fairness is inspired by this construct and corresponds to a great extent to Loi et al. [52]’s concept of design publicity. In line with our idea of broad XAI, Loi et al. [52] demand a form of transparency that explains not only a model and its underlying functioning but also the goals and values that went into its design and how these are embedded in the model. A similar distinction is made by Walmsley [86], who differentiates between functional transparency concerned with the inner workings of an AI model and outward transparency related to communication with stakeholders.
@@
The Role of XAI. In the case of informational fairness, XAI is not an aid to but the desideratum itself. Following this idea, we provide examples of what could be made transparent and communicated to affected parties for each fairness desideratum. Regarding fairness understanding, developers and deployers could justify their fairness objectives, delineate potential tradeoffs, and elucidate the mechanisms through which AI aims to achieve them [52]. For data fairness, they could explain how the training data is composed, highlighting potential sources of bias [88]. Explaining the training objectives and outlining potential uncertainties or goal conflicts might be appropriate to address formal fairness [49]. Regarding perceived fairness, explaining the functions and limitations of XAI-generated information could enhance understanding to approach “appropriate fairness perceptions” [106]. Transparency about the role of AI and the responsibilities of a human decision-maker within the decision-making context might be desirable for fairness with human oversight [139]. The idea of empowering fairness relies on truthful, understandable, and actionable information [81]. Lastly, regarding long-term fairness, unexpected downstream impacts and the factors driving potentially unfair dynamics can be communicated to stakeholders [121].
4 The COMPAS Case
@@
To illustrate the utility of our mapping, we describe how XAI could help address fairness throughout the lifecycle of a high-stakes AI system. As an example, we consider the recidivism prediction software COMPAS developed by Northpointe (today rebranded to equivant Supervision). While COMPAS is not (only) a traditional AI model, it serves as a fitting example because it is data-driven at least to some extent and a blackbox system due its proprietary status [47]. COMPAS has been installed in many US-American jurisdictions in order to predict whether a defendant will likely commit another crime in the near future. Judges use this information, e.g., for decisions about who they release on bail. However, COMPAS has been criticized by investigative journalists at ProPublica for disadvantaging Black people [9]. This is especially problematic given the history of systematic discrimination and marginalization of Black people in the US [140]. In this section, we illustrate both how Northpointe could have addressed different fairness desiderata during system development and how they could still, given COMPAS’ continued use, address some of them. Of course, XAI is not the only means to this end and is not to be conceived as an ethical panacea. Furthermore, our illustration also presupposes an inherent motivation to actually address fairness desiderata, which is not necessarily given in the context of a profit-oriented proprietary system like COMPAS.
@@
Let us proceed along the AI lifecycle. First, Northpointe could have used XAI to ensure that the development is based on an appropriate fairness understanding. ProPublica’s analysis of COMPAS shows that although it was tested for some fairness objective (viz., predictive parity), it falls short on other fairness objectives (viz., equalized odds) in problematic ways [54]. Thus, predictive parity may have been an inadequate fairness objective. Northpointe could also have based its development on insights about the predictive value and correlations of certain features. It has been shown that defendants’ age and number of previous crimes are most predictive for recidivism [141, 47] but also highly correlated to race. Based on such insights, Northpointe could have taken a position on how to treat such features considering pre-existing systematic discrimination.
@@
To address data fairness, Northpointe would have needed to ensure that the data set adequately represents demographic groups targeted by the system and to be aware of existing structural relationships such as statistically higher crime rates in predominantly Black neighborhoods [19]. Descriptive statistics could have already pointed them towards biases introduced in the data collection process—a reason to reiterate the data collection stage to apply new data collection strategies [67]. Statistical analysis could also have helped unveil traces of systematic discrimination in the data (e.g. that Black people are more likely to be arrested for minor offenses due to increased police presence in Black neighborhoods). Northpointe could have used such insights either to change their data collection strategies (e.g., collecting features that are less correlated to race), or they could have used them during the feature selection phase (e.g., to mitigate systematic discrimination at the data-level [66]).
@@
Regarding formal fairness, Northpointe tested COMPAS for equality of error rates (viz., predictive parity). By contrast, journalists at ProPublica tested for the balance of true and false positives (viz., equalized odds) [54]. Notably, many formal fairness metrics cannot be satisfied simultaneously [27], and selecting appropriate formal fairness metrics is an intricate endeavor [3]. However, had Northpointe tested COMPAS for equalized odds, too, they might have anticipated the social backlash and prevented some of the harmful impact. Today, developers can draw on an extensive suite of fairness testing tools for a comprehensive formal fairness assessment [142, 97, 6].
@@
Northpointe could have also detected flaws by engaging with stakeholders and gathering feedback on perceived fairness. Outcome statistics, fairness metrics, or model explanations (e.g., in the form of counterfactuals) could have been presented to a focus group of diverse stakeholders to assess and discuss fairness perceptions (see [109]). Potential concerns could be handled by reconciling the different perspectives, and even today (after deployment) Northpointe could continue to monitor fairness perceptions by providing feedback channels.
@@
Our discussion thus far has primarily focused on measures that COMPAS’ developers could have taken early in the AI lifecycle. But even now, during COMPAS’ continued usage, Northpointe could still address several fairness issues via XAI. First, they could seek fairness with human oversight. This is a particularly relevant desideratum because COMPAS does not make autonomous decisions but informs the decisions of judges. So, Northpointe could (if not already happening) familiarize judges with the basic functioning, strengths, limitations, and uncertainties of the system as well as the factors driving the risk score to ensure judges will use it appropriately. Further, analyzing or supervising the judges’ interactions with the system and its explanations might help to achieve fairness objectives more effectively [74].
@@
Northpointe could also foster empowering fairness by providing defendants with two kinds of explanations: First, global explanations [72] about the general functioning of COMPAS could be communicated. Where relevant factors are under a person’s control (as, e.g., with defendants’ history of misdemeanor or substance abuse), insights about their impact can allow people to make life choices decreasing their risk scores. Second, local explanations [72] could provide information about an individual’s risk score granting the opportunity to (justifiably) contest the score, e.g., by denouncing discriminatory treatment [118], or to (effectively) seek recourse, e.g., by signaling plans to go into rehab [79].
@@
Moreover, Northpointe could monitor long-term fairness by tracking the specified formal fairness objective(s) over time and accordingly adjust the model over several iterations of the AI lifecycle [83]. Further, changes in defendants’ behavior due to specified fairness objectives or increased transparency could be examined. For example, precise global or counterfactual explanations might allow defendants to game the system in unexpected ways strategically [136], which might—similar to unfair recourse [100]—be easier for some than for others.
@@
Finally, Northpointe could have aimed and still could aim, to ensure informational fairness. To this end, they would have needed to document and communicate fairness-related information accordingly. For example, Northpointe could have used model (fairness) cards similar to [49], which would have clarified many of the issues that were only revealed due to the investigative work of Angwin et al. [9]. Most importantly, perhaps, Northpointe could have been candid about the underlying business and fairness objectives from the start, ideally making transparent decisions about tradeoffs (e.g., when looking at predictive parity rather than equalized odds) and justifying why an AI system was an appropriate tool for criminal risk prediction in the first place. In communication with affected parties, COMPAS could be complemented with truthful, understandable, and relevant information [80] explaining the underlying logic. At the same time, jurisdictions could be more transparent about how exactly COMPAS is employed and how judges interact with the system. Overall, the strategic use of XAI along the AI lifecycle could benefit all fairness desiderata regarding a model like COMPAS—given sincere intentions to actually pursue these desiderata.
5 Conclusion and Outlook
@@
We distilled eight fairness desiderata from interdisciplinary literature, mapped them onto the AI lifecycle, and discussed how XAI might be leveraged to contribute to fulfilling them. Finally, we illustrated the utility of our approach by applying it to the COMPAS case. The overall picture we paint (see Figure 2) highlights that to design and utilize XAI for fairness effectively, it is paramount to reflect on which fairness desideratum one seeks to fulfill. Before closing, we shall briefly comment on some limitations of our proposal and point to avenues for future work.
@@
Generalizability. Our discussion has focused primarily on high-risk applications and more traditional decision-support systems. To what extent does our proposal generalize to low-risk applications? And what about generative AI? Consider Google’s multi-modal AI-chatbot Gemini, which has recently received bad publicity for its inaccurate historical depictions—which were driven by misdirected fairness interventions [143]. We suspect that the usage of (seemingly low-risk) AI systems at scale will accumulate to significant societal impacts bearing more subtle threats to fairness such as, e.g., representational harms [89] in the Gemini case. Thus, although contemporary regulation (like the European AI Act) focuses predominantly on high-risk systems, fairness is an important concern for low-risk applications, too. Beyond that, the Gemini case illustrates that fairness considerations affect generative AI just as much as more traditional AI systems for decision-making. Given the rapid advancement and broad adoption of generative AI, we believe that fairness considerations will be indispensable in this context, though they only start to gain traction [144, 145]. What makes them particularly challenging is that traditional stakeholder roles may no longer hold: in the Gemini case, users and affected parties are the same people; and ChatGPT users creating personalized chatbots are no longer “just users” either. Thus, to ensure fairness, future research might not only have to refine fairness desiderata but also rethink stakeholder categories.
Actionability. We acknowledge that our proposal does not present an actionable process model for fair software engineering; nor does it provide regulatory guidelines to be enforced by legal institutions. Still, it may serve as a useful roadmap for researchers, developers, and regulators. For researchers, it provides a starting point for a truly interdisciplinary discourse going beyond discussions of algorithmic fairness focusing on technical aspects of AI systems. For developers, our mapping provides urgently needed guidance as to what fairness challenges should be addressed when developing specific guidelines and requirements (see Hacker [146]). More specifically, it may help determine under which circumstances human oversight is beneficial [74, 43] or how responsibilities can be attributed along the AI lifecycle [147]. Naturally, the implications of our approach will vary with the precise settings; so fine-tuning may be needed for different business, societal, and legal contexts.